The sailboat sails to the northeast at an angle of 65 ° to the east. After 1.2 km and a short stop, the sailboat changes direction and continues to sail to the southeast at an angle of 15 ° to the east, making a shift of 1.8 km. Find the total displacement of the sailboat.


1
Expert's answer
2021-03-02T18:06:28-0500

Let's find "x" and "y" components of the total displacement of the sailboat:


"d_x=1.2\ km\cdot cos65^{\circ}+1.8\ km\cdot cos(360^{\circ}-15^{\circ})=2.24\ km,""d_y=1.2\ km\cdot sin65^{\circ}+1.8\ km\cdot sin(360^{\circ}-15^{\circ})=0.62\ km."

Then, the total displacement can be found from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(2.24\ km)^2+(0.62\ km)^2}=2.32\ km."

We can find the angle as follows:


"\theta=sin^{-1}(\dfrac{d_y}{d})=sin^{-1}(\dfrac{0.62\ km}{2.32\ km})=15.5^{\circ}."


The sign plus means that total displacement has direction "15.5^{\circ} N\ of\ E".

Therefore, the total displacement "d" has magnitude 2.32 km and direction of "15.5^{\circ} N\ of\ E".